OpenStudy (anonymous):

Find 4 consecutive odd integers where the product of the two smaller numbers is 64 less than the product of the two larger numbers. @alli14344 @mathstudent55

OpenStudy (johnweldon1993):

Well I just finished it on your "now closed" question...go have a look And I cannot just give you answers

OpenStudy (johnweldon1993):

Did it at least make sense this time @shortie212 ?

OpenStudy (anonymous):

A way to guarantee that a number will be odd is to call it 2n+1 Odd integers skip every other number, so I can call them 2n+1, 2n+3, 2n+5, 2n+7 given: (2n+1)*(2n+3) = (2n+5)*(2n+7)-64 4n^2+8n+3=4n^2+24n+35-64 8n+3=24n-29 16n=32 n=2

OpenStudy (anonymous):

one sec..

OpenStudy (anonymous):

and ... 2n+1=5 2n+3=7 2n+5=9 2n+7=11 the numbers are 5, 7, 9 , and 11 ..... CHECK: 5*7=9*11-64 35=99-64 35=35 ok :)

OpenStudy (anonymous):

k